OpenStudy (omgthatgirl):

which of the following is the correct notation for sqrt-98+71 a. 7i sqrt 2+71 b. 71+7i sqrt2 c. -7i sqrt 2+71 d. 7 sqr 2+71i

OpenStudy (jhannybean):

Is it \(\sqrt{-98}+71\) or \(\sqrt{-98+71}\)?

OpenStudy (omgthatgirl):

first one @Jhannybean

OpenStudy (omgthatgirl):

@Jhannybean

OpenStudy (jhannybean):

\[\sqrt{-98} \implies \color{red}{\sqrt{-1}}\cdot \sqrt{98}\]\[\color{red}{i=\sqrt{-1}}\]So substitute i for the square root of -1 and add +71

OpenStudy (omgthatgirl):

ok..

OpenStudy (omgthatgirl):

@Jhannybean ?